Assume you have £515 and can exchange CAD908 for £515. Also assume you can exchange $1 for CAD1.0742 and £1 for $1.59. How much arbitrage profit in pounds can you earn? (Ignore transaction costs.)

£15.26

£15.73

£16.07

£16.62

£16.39
Answer #1

Amount of CAD from 515 Pounds =908
Amount of Dollars from 908 CAD =908/1.0742 =845.2802
Amount of Pounds from 845.2802 dollars =845.2802/1.59 =531.62

Profit =531.62-515 =16.62(Option d is correct option)